# A Survey of Behavior Foundation Model: Next-Generation Whole-Body Control System of Humanoid Robots

### v4

@@ V-A5 Multi-agent System
A multi-agent system (MAS) consists of multiple interacting agents that can be either cooperative, competitive, or a mix of both, aiming to address complex tasks that require collaboration, decision-making, and behavior coordination among agents [arai2002advances, gautam2012review, dorri2018multi, madridano2021trajectory, zhou2021survey, ju2022review]. BFMs can fundamentally accelerate the construction of MAS consists of humanoid robots, eliminating the need to effortlessly teach each robot basic survival skills like balance and locomotion before they can collaborate. Instead, researchers can focus directly on higher-level coordination challenges, such as role allocation and team strategy. However, current BFMs trained on single-robot data lack the specialized interaction capabilities needed for optimal collaboration. This presents a promising research direction: developing next-generation BFMs trained explicitly on multi-robot interaction scenarios. Such models could better handle physical coordination challenges, such as object handovers, formation maintenance, and collision avoidance, while preserving their generalizability.
@@ V-A6 Evaluation Mechanism
While foundation models like LLMs benefit from well-established benchmarks (e.g., GPQA [rein2024gpqa] for broad knowledge recall, MATH [hendrycks2measuring] for mathematical problem solving, or MUSR [spraguemusr] for multi-step reasoning), there is no specific and comprehensive evaluation mechanism for evaluating BFM’s capability and guiding the evolution direction [xu2024lvlm, huang2025causality]. A robust assessment of BFMs must consider multiple interdependent factors, including task generalization across unseen scenarios, adaptability to new skills with minimal data, robustness against physical perturbations, and alignment with human safety and interpretability standards. For example, Motivo [tirinzoni2025zeroshot] combines quantitative metrics, such as task success rates, with qualitative human evaluations of motion naturalness; yet, critical gaps remain in assessing compositional skill combinations, hardware-specific constraints, and long-term behavioral stability. Future benchmarks should focus on progressive difficulty levels and cross-domain transfer tests to effectively assess the potential of BFMs as general-purpose physical controllers in both simulated and real-world scenarios. This approach will ultimately guide the field toward developing more capable and reliable humanoid systems.
@@ V-B1 Ethical Issues
Ethical issues consistently accompany the development of diverse foundation models [weidinger2021ethical, yan2024practical, bommasani2024considerations], which involve biased or unlicensed data, racial discrimination, and uncontrollable behaviors, etc. For BFMs, training on non-diverse motion datasets may encode demographic biases such as favoring movements natural to specific age groups or body types, which then propagate into robotic behaviors and create embodied forms of discrimination. Meanwhile, privacy risks escalate beyond data memorization to movement analytics, where rehabilitation or performance data could leak sensitive health information through generated motions. The physical instantiation of BFMs introduces unprecedented risks: unlike purely digital models, misaligned BFMs might reproduce unsafe or socially harmful behaviors (e.g., aggressive gestures, or exclusionary motions) with real-world consequences. While techniques like differential privacy and federated learning offer partial solutions, they struggle with the temporal nature of continuous motion data. In summary, BFMs demand novel governance frameworks that address both data provenance and normativity of real-time behavior.
@@ V-B2 Safety Mechanism
As BFMs will be increasingly deployed in real-world robotic systems, they introduce critical safety requirements beyond those of digital foundation models [vasic2013safety, lasota2017survey, kumagai2019toward]. A key issue is maintaining model behavior integrity, particularly in safety-critical scenarios such as human-robot interaction and autonomous navigation. When trained on large-scale yet weakly curated motion datasets, BFMs may unintentionally learn unsafe or undesirable behaviors. Even minor changes in sensory input, caused by adversarial attacks or sensor noise, can lead to control failures. This highlights the need for robustness against shifts in data distribution and protection against malicious input manipulation.
While most current BFMs focus primarily on proprioceptive inputs, integrating multimodal information (e.g., visual, linguistic, and auditory cues) has emerged as a promising direction for more generalizable and situationally aware control. However, multimodality introduces new vulnerabilities. Adversaries can exploit inconsistencies across modalities, as seen in the well-known CLIP case where an apple image was misclassified as an ”iPod” due to an overlaid label [goh2021multimodal]. Such cross-modal confusion can be especially dangerous when a BFM is adapted for an unimodal task but retains sensitivity to irrelevant signals from other modalities. These challenges underscore the need to develop robust safety mechanisms for BFMs. Future work could prioritize adversarial robustness, cross-modal consistency checks, and disentangling modality-specific information to ensure predictable, trustworthy robot behavior in open-world environments.

### v5

@@ V-A5 Multi-agent System
A multi-agent system (MAS) consists of multiple interacting agents that can be either cooperative, competitive, or a mix of both, aiming to address complex tasks that require collaboration, decision-making, and behavior coordination among agents [205, 206, 207, 208, 209, 210]. BFMs can fundamentally accelerate the construction of MAS consists of humanoid robots, eliminating the need to effortlessly teach each robot basic survival skills like balance and locomotion before they can collaborate. Instead, researchers can focus directly on higher-level coordination challenges, such as role allocation and team strategy. However, current BFMs trained on single-robot data lack the specialized interaction capabilities needed for optimal collaboration. This presents a promising research direction: developing next-generation BFMs trained explicitly on multi-robot interaction scenarios. Such models could better handle physical coordination challenges, such as object handovers, formation maintenance, and collision avoidance, while preserving their generalizability.
@@ V-A6 Evaluation Mechanism
While foundation models like LLMs benefit from well-established benchmarks (e.g., GPQA [211] for broad knowledge recall, MATH [212] for mathematical problem solving, or MUSR [213] for multi-step reasoning), there is no specific and comprehensive evaluation mechanism for evaluating BFM’s capability and guiding the evolution direction [214, 215]. A robust assessment of BFMs must consider multiple interdependent factors, including task generalization across unseen scenarios, adaptability to new skills with minimal data, robustness against physical perturbations, and alignment with human safety and interpretability standards. For example, Motivo [53] combines quantitative metrics, such as task success rates, with qualitative human evaluations of motion naturalness; yet, critical gaps remain in assessing compositional skill combinations, hardware-specific constraints, and long-term behavioral stability. Future benchmarks should focus on progressive difficulty levels and cross-domain transfer tests to effectively assess the potential of BFMs as general-purpose physical controllers in both simulated and real-world scenarios. This approach will ultimately guide the field toward developing more capable and reliable humanoid systems.
@@ V-B1 Ethical Issues
Ethical issues consistently accompany the development of diverse foundation models [216, 217, 218], which involve biased or unlicensed data, racial discrimination, and uncontrollable behaviors, etc. For BFMs, training on non-diverse motion datasets may encode demographic biases such as favoring movements natural to specific age groups or body types, which then propagate into robotic behaviors and create embodied forms of discrimination. Meanwhile, privacy risks escalate beyond data memorization to movement analytics, where rehabilitation or performance data could leak sensitive health information through generated motions. The physical instantiation of BFMs introduces unprecedented risks: unlike purely digital models, misaligned BFMs might reproduce unsafe or socially harmful behaviors (e.g., aggressive gestures, or exclusionary motions) with real-world consequences. While techniques like differential privacy and federated learning offer partial solutions, they struggle with the temporal nature of continuous motion data. In summary, BFMs demand novel governance frameworks that address both data provenance and normativity of real-time behavior.
@@ V-B2 Safety Mechanism
As BFMs will be increasingly deployed in real-world robotic systems, they introduce critical safety requirements beyond those of digital foundation models [219, 220, 221]. A key issue is maintaining model behavior integrity, particularly in safety-critical scenarios such as human-robot interaction and autonomous navigation. When trained on large-scale yet weakly curated motion datasets, BFMs may unintentionally learn unsafe or undesirable behaviors. Even minor changes in sensory input, caused by adversarial attacks or sensor noise, can lead to control failures. This highlights the need for robustness against shifts in data distribution and protection against malicious input manipulation.
While most current BFMs focus primarily on proprioceptive inputs, integrating multimodal information (e.g., visual, linguistic, and auditory cues) has emerged as a promising direction for more generalizable and situationally aware control. However, multimodality introduces new vulnerabilities. Adversaries can exploit inconsistencies across modalities, as seen in the well-known CLIP case where an apple image was misclassified as an ”iPod” due to an overlaid label [222]. Such cross-modal confusion can be especially dangerous when a BFM is adapted for an unimodal task but retains sensitivity to irrelevant signals from other modalities. These challenges underscore the need to develop robust safety mechanisms for BFMs. Future work could prioritize adversarial robustness, cross-modal consistency checks, and disentangling modality-specific information to ensure predictable, trustworthy robot behavior in open-world environments.
@@ VI Conclusion
Acknowledgments
This work is supported, in part, by the Hong Kong SAR Research Grants Council under Grant No. PolyU 15224823, the Guangdong Basic and Applied Basic Research Foundation under Grant No. 2024A1515011524, the NSFC under Grant No. 62302246, the ZJNSFC under Grant No. LQ23F010008, the Ningbo under Grants No. 2023Z237 & 2023CX050011 & 2024Z284 & 2024Z289 & 2025Z038, and the Ningbo Institute of Digital Twin (IDT) under Grant No. S203.2.01.32.002. This work is also supported by the High Performance Computing Center at the IDT and the Eastern Institute of Technology, Ningbo.